A skateboarder, starting from rest, rolls down a 13.0-m ramp. When she arrives at the bottom of the ramp her speed is 8.40 m/s.

(a) Determine the magnitude of her acceleration, assumed to be constant.


(b) If the ramp is inclined at 25.5° with respect to the ground, what is the component of her acceleration that is parallel to the ground?
Answer #1

(a) The magnitude of her acceleration, assumed to be constant which will be given as :

using equation of motion 3,   vf2 = v02 + 2 a x                                                                { eq.1 }

where, v0 = initial speed of skateboarder = 0 m/s

x = distance travelled = 13 m

vf = final speed at the bottom of ramp = 8.4 m/s

inserting all these values in above eq.

(8.4 m/s)2 = (0 m/s)2 + 2 a (13 m)

(70.56 m2/s2) = (26 m) a

a = (70.56 m2/s2) / (26 m)

a = 2.71 m/s2

(b) If the ramp is inclined at 25.50 with respect to the ground, the component of her acceleration that is parallel to the ground which is given as ::

ax = a cos \theta                                                                  { eq.2 }

inserting the values in eq.2,

ax = (2.71 m/s2) cos (25.50)

ax = 2.44 m/s2